Grade 12Algebra

The value of sigma r=1 to 10 r*nCr/nCr-1

Brown sam
12 Years agoGrade 12

3 Answers

Arun Kumar
ApprovedApproved Tutor Answer12 Years ago
Hello Student,
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
S Sharma
7 Years ago
Sir,
The 4th step should be..
10n-45
As we can see that in the last step
There is 5(2n-9)
$o please look in the issue 
If I'm correct or not....
       Well thank you for the answer sir
I'm very well impressed with you...
 
 
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem.
 
We know, nCr/nCr-1 = (n – r + 1)/r
Hence, ∑ r. nCr/nCr-1 = ∑ (n – r + 1)      [from r = 1 to r = 10]
= ∑ (n + 1)  –  ∑ r
= 10(n + 1) – 10(10 + 1)/2
= 10n + 10 – 55
= 10n – 45
= 5(2n – 9)
 
Thanks and regards,
Kushagra